Jebediah:

Two lines, C and D, are represented by the equations given below: Line C: y = x + 10 Line D: y = 3x + 2 Which of the following shows the solution to the system of equations and explains why? (5 points) Please post all aspects of how you got there so I know for next time :P     (5, 17), because one of the lines passes through this point (5, 17), because the point lies between the two axes (4, 14), because the point does not lie on any axis (4, 14), because both lines pass through this point

Jebediah:

@dude

dude:

In order to find the solution you must set them equal to each other \(x+10 = 3x + 2\\ x=4\) Substitute to find y \(4+10=14\\ (4,14)\) A solution is the point where both lines intersect

Jebediah:

THANKS SO MUCH